An external voltage of 60 V is applied to a coil whose resistance is 5.25 Ω. What is the current in the coil if the back EMF is 9.85 Ω?




1
Expert's answer
2021-12-09T09:56:02-0500

"\u03f5i=V_S\u2212IR"

"I=-[\frac{ϵi-V_S}{R}]"


Where;

ϵi=back EMF=9.35V

"V_s" =External voltage=60V

I= current=

R=resistance=5.25Ω


"I=-[\frac{9.85-60}{5.25}]=9.65A"

I=9.65A